OpenStudy (anonymous):

Match the following items by evaluating the expression for x = -2. x -2^ x -1^ x 0^ x 1^ x 2^

OpenStudy (solomonzelman):

x^0 is 1 for any x (even when x=0, on some level). So (-2)^0=0 (-2)^(-2)= 1/(-2)²=1/4 (-2)^(-1)=1/(-2)¹ = -1/2 (-2)¹ = -2 (-2)² = 4

OpenStudy (solomonzelman):

Don't really get the word match in your question.